# Robots that Suggest Safe Alternatives

### v1

@@ I Introduction
In this paper, we want to endow robots with the ability to know when they can confidently execute a user’s desired goal and propose safe alternatives when they cannot. Specifically, we study goal-conditioned robot policies [1] such as those obtained via imitation learning [2]. While this paradigm has enabled robots to learn complex behaviors and adapt to specified goals online [3, 4], these learned policies can degrade when faced with out-of-distribution goal requests or states [5]. In other words, given a pre-trained goal-conditioned policy, it is hard to ensure that the closed-loop robot behavior will always be safe (e.g., doesn’t collide with the environment) and performant (e.g., will successfully pick up the cup) for any new user goal and initial state. Prior works have quantified policy uncertainty [6] or developed out-of-distribution input detectors [7], but these approaches are only a monitoring mechanism; they don’t provide a way for the robot to actively propose alternatives that it can accomplish safely and effectively.
@@
To close this gap, we propose that robots suggest safe alternatives. Our approach is inspired by safety filtering techniques from control theory [8, 9]. Traditional safety filters take a candidate robot action (e.g., generated by a pre-trained policy) and minimally adjust (i.e., “filter”) it to be safe. The filtering mechanism projects the candidate action onto the safe and live (i.e. goal-reaching) control set, where the control set is computed via methods like Control Barrier [10] or Control Lyapunov Functions [11, 12], Hamilton-Jacobi (HJ) reachability analysis [13, 14], or model predictive filters [15]. However, this action-space filtering does not enable the robot to naturally suggest alternatives in a human-verifable way. Our key idea is that an
@@
We leverage HJ reachability analysis to synthesize a goal-conditioned reach-avoid value function that is used within our safe control framework. This computation is performed once offline and automatically quantifies how capable the robot’s pre-trained policy is at accomplishing the task while staying safe, for a suite of possible goal inputs. Due to the dimensionality of the problem induced by goal parameterization, we leverage principled but approximate reachability solvers rooted in reinforcement learning [16], and empirically quantify the error of our learned value function. Online, we use this reach-avoid value function to monitor if a specified goal meets the reach-avoid criteria and perform safety filtering over the goals to suggest alternatives if not.
@@ II Related Work
Safety Filtering. Safety filters—which detect unsafe actions and minimally modify them—are increasingly popular ways to ensure closed-loop safety [17, 18, 19, 9]. The most popular methods are control barrier functions (CBFs) [10, 20, 21, 22, 23], Hamilton-Jacobi (HJ) reachability [24, 25, 26, 27, 28], and model predictive shielding [29]. In this work we build off of HJ reachability due to its ability to handle non-convex target and constraint sets, control constraints and uncertainty in the system dynamics, and its association with a suite of numerical tools including recent neural approximations that scaled safe set synthesis to 15-200 dimensions [30, 31, 32, 33]. Our key idea is that by treating the human’s goal as a virtual state, we can do safety value function synthesis and safety filtering on the goal (instead of on the actions as is typical). This enables the robot to minimally modify the human’s desired goal and propose safe alternatives.
Uncertainty Quantification of Learned Robot Policies. For modular robot policies that utilize an upstream goal or intent estimator, prior works have quantified goal uncertainty [34, 35, 36, 37], calibrated task plans inferred from language commands [38] and quantified their execution risk [39]. For end-to-end behavior cloned policies, prior works have quantified their generalizability via statistical bounds [40], action uncertainty via temperature scaling [41] or conformal prediction [42], and predicted policy success rate via value estimation [43]. Our work uses control-theoretic verification tools to analyze the closed-loop success of a robot’s policy.
Robot Communication of Uncertainty & Capability. Prior works in human-robot interaction have enabled robots to communicate their task uncertainty via dialogue [44, 45, 46], communicate their objectives through motion or haptics [47, 48], express physical capabilities [49], or explain their failures [50, 51] to people (see [52] for a review). Instead of having robots only explain what they are uncertain about (or ask for help), we enable robots to actively suggest alternatives they can safely accomplish.
@@ III Method: Suggesting Safe Alternatives
Offline: Reach-Avoid Analysis of Goal-Conditioned Robot Policies. We use reachability analysis to automatically verify for which initial states and desired goals can the robot’s policy satisfy both safety and liveness constraints. Our core idea is to treat the space of possible goals that the human could input ($g\in\mathcal{G}$) as a virtual state that has zero dynamics (i.e., $g_{t+1}=g_{t}$) during the reachability analysis. This enables our approach to quantify the sensitivity of the robot’s policy to the range of possible goal inputs.
For offline analysis, we leverage Hamilton-Jacobi (HJ) reachability [14, 53]. This verification technique is compatible with nonlinear dynamical systems, arbitrary non-convex failure and target set representations, and has a suite of associated numerical tools, from exact grid-based solvers [54] and conservative linearization [55] to approximate but scalable techniques such as reinforcement learning [30, 16] and self-supervised learning [31]. We encode our safety specification via a failure set $\mathcal{F}\subset\mathcal{S}$ (e.g., the object slipped from the gripper) and liveness via a target set $\mathcal{T}\subset\mathcal{S}$ (e.g., the object height must be above the table).
For computation, the target and failure sets are encoded via Lipschitz continuous margin functions $l(\cdot)$ and $h(\cdot)$ respectively: $\mathcal{T}:=\{s\mid l(s)\leq 0\}$ and :=$\mathcal{F}:=\{s\mid h(s)>0\}$. One such function is the signed distance to the set boundary. Intuitively, these margin functions will measure the “closest” the robot’s policy $\pi(s;g)$ ever got to violating safety and accomplishing the task given the specified goal $g$. We introduce goal-conditioned target and failure sets, $\mathcal{T}_{g}$ and $\mathcal{F}_{g}$, as well as corresponding goal-conditioned margin function, $l_{g}(s)$ and $h_{g}(s)$. With these in hand, we can define the goal- and policy-conditioned safety value function as:
$V_{*}^{\pi}(s;g)=$ $\max\min_{\tau\in\{0,1,\ldots\}}\Big{\{}l_{g}(\xi^{\pi(\cdot;g)}_{s}(\tau)),$
@@
Where the robot’s trajectory starting from state $s$ and applying policy $\pi(\cdot;g)$ is denoted by $\xi^{\pi(\cdot;g)}_{s}$. If $\xi^{\pi(\cdot;g)}_{s}$ violates $\mathcal{F}$ at any time, then the inner maximum will be positive and thus the overall value will be positive as well. If the robot’s trajectory never enters $\mathcal{F}$, then the overall value will be negative if and only if the robot reaches the target $\mathcal{T}$ (which is encoded via the sub-zero level set of $l$). The outer minimum ensures that the value function “remembers” these events over the entire time horizon. In other words, the value can only be negative if the target $\mathcal{T}$ is reached without ever violating the safety constraint $\mathcal{F}$ along the way. Following prior work [16], it can be shown that this value function must satisfy the fixed-point reach-avoid Bellman equation:
@@
where $s^{\pi}_{+}:=f(s,\pi(s;g))$ is the next state the robot reaches after applying the control from the goal-conditioned policy.
To scale reach-avoid analysis to higher dimensional state and goal spaces, we leverage the principled modification of Eq. 2 introduced in [16], rendering the reach-avoid problem compatible with reinforcement learning approximations:
@@
$g_{\mathcal{R}}=\arg\min_{g\in\mathcal{G}}~{}d(g,g_{\mathcal{H}})$ (4)
@@
where $d(g,g_{\mathcal{H}})$ is a distance metric. If the human’s original goal $g_{\mathcal{H}}$ and robot’s initial state $s$ does not satisfy safety and liveness (i.e., $V^{\pi}(s;g_{\mathcal{H}})>0$), then the optimization above will be solved to find an alternative goal $g_{\mathcal{R}}$ that’s similar to $g_{\mathcal{H}}$. In our experiments, we use the Euclidian distance $d(g,g_{\mathcal{H}})$ between the goal representations.
@@ IV Experimental Setup
Base Robot Policy ($\pi$). For all environments, the robot’s base policy is trained via behavior cloning (BC). For navigation, expert demonstrations $(s,a,g)\sim{\mathcal{D}}$ are collected by using an optimal control law obtained from a reach-avoid computation obtained via the optimizedDP solver [57]. We pre-compute three separate optimal controllers for three goal locations, $g_{\mathcal{D}}\in\{(3,-3),(3,0),(3,3)\}$ and obtain 50 successful expert trajectories per goal. The BC policy is a 3-layer MLP with 128 neurons per layer and RELU nonlinearity and trained via the mean squared error loss $\mathcal{L}_{BC}(\mathcal{D})=\mathbb{E}_{(s^{j},a^{j},g^{j})\sim\mathcal{D}}\left\|\pi(s^{j};g^{j})-a^{j}\right\|_{2}^{2}$, where $j$ is each state-goal-action pair. We train for 500 epochs with a 100 epoch early stop patience. In manipulation, we collected demonstrations in the robomimic [58] environment via expert human teleoperation, and obtained 100 successful demonstrations per object. The BC policy is a 2-layer MLP with 1024 neurons per layer and ReLU nonlinearity. Its tra ned with the AdamW [59] optimizer and for 100 epochs. Every 5 epochs, we rollout the policy 50 times, and use the policy with the highest success rate.
@@ V Experimental Results
To understand each component of SALT, we study three questions in simulation experiments: (1) how accurate is the reach-avoid value function approximation?, (2) how does SALT compare to alternative runtime monitoring schemes (e.g., ensembles)?, and (3) how aligned are the alternatives that SALT proposes with human-acceptable ones?
@@ V-A How Accurate is SALT’s Reach-Avoid Value Function?
Metrics. To evaluate the reliability of the learned value network, we measure the true success rate (TSR: network predicts the policy can safely accomplish the task and in reality it can), true failure rate (TFR: network predicts the policy cannot safely accomplish the task and in reality it cannot), false success rate (FSR: network predicts the policy can safely accomplish the task but in reality it cannot), and false failure rate (FFR: network predicts the policy cannot safely accomplish the task but in reality in can).
Evaluation Approach. For the low-dimensional (4D) navigation example, we evaluate our approximation for a dense gridding of the state space (50x50x20x20). For the high-dimensional manipulation example, exhaustive gridding is not feasible. Thus, we randomly sample initial physical states $s$ and goals $g$, perform a rollout in our simulator, and check if the value of the network accurately reflects the outcome of policy execution (a safe success, or not). We do this on 1,000 initial $(s,g)$ pairs sampled near the zero level set of the approximate value network ($V^{\pi}(s,g)\approx 0$), since a accurate boundary matters for monitoring, and across 10 random seeds.
@@
Environment TSR (%) TFR (%) FSR (%) FFR (%)
Navigation 55.51 39.48 1.28 3.72
Manipulation 59.98 ($\pm$2.10) 14.55 ($\pm$1.43) 12.64 ($\pm$1.28) 12.83 ($\pm$1.13)
@@ V-B What is the Benefit of SALT as a Runtime Monitor?
Method TNR % ($\uparrow$) TPR % ($\uparrow$) FPR % ($\downarrow$) FNR % ($\downarrow$) TNR % ($\uparrow$) TPR % ($\uparrow$) FPR % ($\downarrow$) FNR % ($\downarrow$)
Ensemble 65.97 ($\pm$0.88) 1.92 ($\pm$0.38) 12.31 ($\pm$0.48) 19.80 ($\pm$0.52) 34.61 ($\pm$1.68) 27.45 ($\pm$1.96) 32.65 ($\pm$1.27) 5.26 ($\pm$0.89)
RewardSum 53.35 ($\pm$1.34) 5.95 ($\pm$0.74) 4.87 ($\pm$0.60) 35.78 ($\pm$1.26) 64.47 ($\pm$1.61) 4.67 ($\pm$0.68) 8.85 ($\pm$1.01) 22.01 ($\pm$1.21)
SALT (ours) 57.00 ($\pm$1.97) 37.34 ($\pm$2.03) 1.28 ($\pm$0.15) 4.38 ($\pm$0.78) 61.21 ($\pm$1.91) 13.23 ($\pm$1.56) 13.23 ($\pm$1.23) 12.33 ($\pm$0.85)

### v2

@@ I Introduction
In this paper, we want to endow robots with the ability to know when they can confidently execute a user’s desired goal and propose safe alternatives when they cannot. Specifically, we study goal-conditioned robot policies [1] such as those obtained via imitation learning [2]. While this paradigm has enabled robots to learn complex behaviors and adapt to specified goals online [3], these learned policies can degrade when faced with out-of-distribution goal requests or states [4]. In other words, given a pre-trained goal-conditioned policy, it is hard to ensure that the closed-loop robot behavior will always be safe (e.g., doesn’t collide with the environment) and performant (e.g., will successfully pick up the cup) for any new user goal and initial state. Prior works have quantified policy uncertainty [5] or developed out-of-distribution input detectors [6], but these approaches are only a monitoring mechanism; they don’t provide a way for the robot to actively propose alternatives that it can accomplish safely and effectively.
@@
To close this gap, we propose that robots suggest safe alternatives. Our approach is inspired by safety filtering techniques from control theory [7]. Traditional safety filters take a candidate robot action (e.g., generated by a pre-trained policy) and minimally adjust (i.e., “filter”) it to be safe. The filtering mechanism projects the candidate action onto the safe and live (i.e. goal-reaching) control set, where the control set is computed via methods like Control Barrier [8] or Control Lyapunov Functions [9], Hamilton-Jacobi (HJ) reachability analysis [10], or model predictive filters [11]. However, this action-space filtering does not enable the robot to naturally suggest alternatives in a human-verifable way. Our key idea is that an
@@
We leverage HJ reachability analysis to synthesize a goal-conditioned reach-avoid value function that is used within our safe control framework. This computation is performed once offline and automatically quantifies how capable the robot’s pre-trained policy is at accomplishing the task while staying safe, for a suite of possible goal inputs. Due to the dimensionality of the problem induced by goal parameterization, we leverage principled but approximate reachability solvers rooted in reinforcement learning [12], and empirically quantify the error of our learned value function. Online, the reach-avoid value function takes in the current state and a specified goal and determines whether the goal meets the reach-avoid criteria (safe and live). If not, we perform safety filtering over the goals to suggest alternatives.
@@ II Related Work
Safety Filtering. Safety filters—which detect unsafe actions and minimally modify them—are increasingly popular ways to ensure closed-loop safety [13, 14]. The most popular methods are control barrier functions (CBFs) [8], Hamilton-Jacobi (HJ) reachability [15, 16], and model predictive shielding [17].
In this work we build off of HJ reachability due to its ability to handle non-convex target and constraint sets, control constraints and uncertainty in the system dynamics, and its association with a suite of numerical tools including recent neural approximations that scaled safe set synthesis to 15-200 dimensions [18, 19]. Our key idea is that by treating the human’s goal as a virtual state, we can do safety value function synthesis and safety filtering on the goal (instead of on the actions as is typical). This enables the robot to minimally modify the human’s desired goal and propose safe alternatives.
Uncertainty Quantification of Learned Robot Policies. For modular robot policies that utilize an upstream goal or intent estimator, prior works have quantified goal uncertainty [20], calibrated task plans inferred from language commands [21] and quantified their execution risk [22]. For end-to-end behavior cloned policies, prior works have quantified their generalizability via statistical bounds [23], action uncertainty via temperature scaling [24], and predicted policy success rate via value estimation [25].
Our work uses control-theoretic verification tools to analyze the closed-loop success of a robot’s policy.
Robot Communication of Uncertainty & Capability. Prior works in human-robot interaction have enabled robots to communicate their task uncertainty via dialogue [26], communicate their objectives through motion or haptics [27, 28], express physical capabilities [29], or explain their failures [30] to people (see [31] for a review). Instead of having robots only explain what they are uncertain about (or ask for help), we enable robots to actively suggest alternatives they can safely accomplish.
@@ III Method: Suggesting Safe Alternatives
Offline: Reach-Avoid Analysis of Goal-Conditioned Robot Policies. We use reachability analysis to automatically verify for which initial states and desired goals can the robot’s policy satisfy both safety and liveness constraints. Our core idea is to treat the space of possible goals that the human could ask about at deployment time ($g\in\mathcal{G}$) as a virtual state that has zero dynamics (i.e., $\dot{g}=0$) during the reachability analysis. This approach, inspired by parameter-conditioned reachability [32], enables us to quantify the safety and liveness sensitivity of the robot’s policy as a function of all possible goal inputs.
For offline analysis, we leverage Hamilton-Jacobi (HJ) reachability. This verification technique is compatible with nonlinear dynamical systems, arbitrary non-convex failure and target set representations, and has a suite of associated numerical tools, from exact grid-based solvers [33] to approximate but scalable techniques such as reinforcement learning [18, 12] and self-supervised learning [19].
We encode our safety specification via a failure set $\mathcal{F}\subset\mathcal{S}$ (e.g., the object slipped from the gripper) and liveness via a target set $\mathcal{T}\subset\mathcal{S}$ (e.g., the object height must be above the table). For computation, the target and failure sets are encoded via Lipschitz continuous margin functions $l(\cdot)$ and $h(\cdot)$ respectively: $\mathcal{T}:=\{s\mid l(s)\leq 0\}$ and :=$\mathcal{F}:=\{s\mid h(s)>0\}$. One such function is the signed distance to the set boundary. Intuitively, these margin functions will measure the “closest” the robot’s policy $\pi(s;g)$ ever got to violating safety and accomplishing the task given the specified goal $g$. We introduce goal-conditioned target and failure sets, $\mathcal{T}_{g}$ and $\mathcal{F}_{g}$, as well as corresponding goal-conditioned margin function, $l_{g}(s)$ and $h_{g}(s)$. With these in hand, we can define the goal- and policy-conditioned safety value function as:
$V_{*}^{\pi}(s;g)=$ $\min_{\tau\in\{0,1,\ldots\}}\max\Big{\{}l_{g}(\xi^{\pi(\cdot;g)}_{s}(\tau)),$
@@
where the robot’s trajectory starting from state $s$ and applying policy $\pi(\cdot;g)$ is denoted by $\xi^{\pi(\cdot;g)}_{s}$. Intuitively, the outer maximum acts as a mechanism to remember if the robot has ever entered the failure set ${\mathcal{F}}$ up to this time (right-hand side) and has satisfied the liveness property (left-hand side). If $\xi^{\pi(\cdot;g)}_{s}$ enters $\mathcal{F}$ at any time, then the inner maximum will be positive, and thus the overall value will also be positive. In contrast, if the robot’s trajectory never enters $\mathcal{F}$, then the overall value will be negative if and only if the robot reaches the target $\mathcal{T}$ (which is encoded via the subzero level set of $l$). The outer minimum ensures that the value function “remembers” these events over the entire time horizon. In other words, the value can only be negative if the target $\mathcal{T}$ is reached without ever violating the safety constraint $\mathcal{F}$ along the way.
Following prior work [12], it can be shown that this value function must satisfy the fixed-point reach-avoid Bellman equation:
@@
where $s^{\pi}_{+}:=f(s,\pi(s;g))$ is the next state the robot reaches after applying the control from the goal-conditioned policy $\pi(\cdot;g)$ and the * indicates optimality of the reach-avoid value function, $V_{*}^{\pi}$, computed with a perfect solver.
To scale reach-avoid analysis to higher dimensional state and goal spaces, we leverage the principled time-discounted formulation of Eq. 2 introduced in [12], rendering the reach-avoid problem compatible with reinforcement learning approximations (e.g., Q-learning, REINFORCE [34, 18]):
@@
$g_{\mathcal{R}}=\arg\min_{g\in\mathcal{G}}~{}d(\mathcal{E}(g),\mathcal{E}(g_{\mathcal{H}});\theta)$ (4)
@@
where $d(\cdot)$ is a similarity measure, and $\mathcal{E}(\cdot)$ is an encoder that maps goal representations to the goal space. The similarity measure is parameterized by human intent $\theta$, since similarity notion differs user by user. We stress that the representation of the goals $g$, and the corresponding similarity measure $d(\cdot)$ are a key design decision, and one which we explore in Sec. V-C. If the human’s original goal $g_{\mathcal{H}}$ and robot’s initial state $s$ does not satisfy safety and liveness (i.e., $V^{\pi}(s;g_{\mathcal{H}})>0$), then the optimization above will be solved to find an alternative goal $g_{\mathcal{R}}$ that’s similar to $g_{\mathcal{H}}$. In our experiments, we first start with pose-based representations of $g$ and an Euclidean distance function (not conditioned on intent), then explore semantic similarity and goal representations (conditioned on intent).
@@ IV Experimental Setup
Base Robot Policy ($\pi$). For all environments, the robot’s base policy is trained via behavior cloning (BC). For navigation, expert demonstrations $(s,a,g)\sim{\mathcal{D}}$ are collected by using an optimal control law obtained from a reach-avoid computation obtained via the optimizedDP solver [36]. We pre-compute three separate optimal controllers for three goal locations, $g_{\mathcal{D}}\in\{(3,-3),(3,0),(3,3)\}$ and obtain 50 successful expert trajectories per goal. The BC policy is a 3-layer MLP with 128 neurons per layer and RELU nonlinearity and trained via the mean squared error loss $\mathcal{L}_{BC}(\mathcal{D})=\mathbb{E}_{(s^{j},a^{j},g^{j})\sim\mathcal{D}}\left\|\pi(s^{j};g^{j})-a^{j}\right\|_{2}^{2}$, where $j$ is each state-goal-action pair. We train for 500 epochs with a 100 epoch early stop patience. In manipulation, we collected demonstrations in the robomimic [37] environment via expert human teleoperation, and obtained 100 successful demonstrations per object. The BC policy is a 2-layer MLP with 1024 neurons per layer and ReLU nonlinearity. It is trained with the AdamW [38] optimizer and for 100 epochs. Every 5 epochs, we rollout the policy 50 times, and use the policy with the highest success rate.
@@ V Experimental Results
To understand each component of SALT, we study four questions in simulation experiments: (1) how accurate is the reach-avoid value function approximation?, (2) how does SALT compare to alternative runtime monitoring schemes (e.g., ensembles)?, (3) how should SALT measure “similar” alternative goals?, and (4) how aligned are the alternatives that SALT proposes with human-acceptable ones?
@@ V-A How Accurate is SALT’s Reach-Avoid Value Function?
Metrics. To evaluate the reliability of the learned value network, we measure the true success rate (TSR: network predicts the policy can safely accomplish the task and in reality it can), true failure rate (TFR: network predicts the policy cannot safely accomplish the task and in reality it cannot), false success rate (FSR: network predicts the policy can safely accomplish the task but in reality it cannot), and false failure rate (FFR: network predicts the policy cannot safely accomplish the task but in reality in can). $F_{1}$-Score represents the predictive performance of a binary classifier (1.0 indicates perfect precision).
Evaluation Approach. For the low-dimensional (4D) navigation example, we evaluate our approximation for a dense gridding of the state space (50x50x20x20). For the high-dimensional manipulation example, exhaustive gridding is not feasible. Thus, we randomly sample initial physical states $s$ and goals $g$, perform a rollout in our simulator, and check if the value of the network accurately reflects policy execution outcome (safe success, or not). We do this on 1,000 initial $(s,g)$ pairs sampled near the zero level set of the approximate value network ($V^{\pi}(s,g)\approx 0$), since a accurate boundary matters for monitoring, and across 10 random seeds.
@@
Environment TSR (%) TFR (%) FSR (%) FFR (%) $F_{1}$-Score
Navigation 55.51 39.48 1.28 3.72 0.96
Manipulation 59.98 ($\pm$2.10) 14.55 ($\pm$1.43) 12.64 ($\pm$1.28) 12.83 ($\pm$1.13) 0.82
@@ V-B What is the Benefit of SALT as a Runtime Monitor?
Method TNR % ($\uparrow$) TPR % ($\uparrow$) FPR % ($\downarrow$) FNR % ($\downarrow$) $F_{1}$-Score ($\uparrow$) TNR % ($\uparrow$) TPR % ($\uparrow$) FPR % ($\downarrow$) FNR % ($\downarrow$) $F_{1}$-Score ($\uparrow$)
Ensemble 65.97 ($\pm$0.88) 1.92 ($\pm$0.38) 12.31 ($\pm$0.48) 19.80 ($\pm$0.52) 0.80 34.61 ($\pm$1.68) 27.45 ($\pm$1.96) 32.65 ($\pm$1.27) 5.26 ($\pm$0.89) 0.65
RewardSum 53.35 ($\pm$1.34) 5.95 ($\pm$0.74) 4.87 ($\pm$0.60) 35.78 ($\pm$1.26) 0.72 64.47 ($\pm$1.61) 4.67 ($\pm$0.68) 8.85 ($\pm$1.01) 22.01 ($\pm$1.21) 0.81
SALT (ours) 57.00 ($\pm$1.97) 37.34 ($\pm$2.03) 1.28 ($\pm$0.15) 4.38 ($\pm$0.78) 0.95 61.21 ($\pm$1.91) 13.23 ($\pm$1.56) 13.23 ($\pm$1.23) 12.33 ($\pm$0.85) 0.83